if log 784 = 2.8943, then log (786 X10⁴) is


1
Expert's answer
2021-05-20T10:08:16-0400

"\log(784\cdot10^4)=\log784+\log10^4="


"=2.8943+4\cdot\log10=2.8943+4\cdot1=2.8943+4=6.8943" . Answer